# A faster FPRAS for #NFA

### 20 Dec 2023 v1

@@ Abstract.
Given a non-deterministic finite automaton (NFA) $\mathcal{A}$ with $m$ states, and a natural number $n\in\mathbb{N}$ (presented in unary), the #NFA problem asks to determine the size of the set $\mathcal{L}(\mathcal{A}_{n})$ of words of length $n$ accepted by $\mathcal{A}$. While the corresponding decision problem of checking the emptiness of $\mathcal{L}(\mathcal{A}_{n})$ is solvable in polynomial time, the #NFA problem is known to be #P-hard. Recently, the long-standing open question — whether there is an FPRAS (fully polynomial time randomized approximation scheme) for #NFA — was resolved in (Arenas et al., 2019). The FPRAS due to (Arenas et al., 2019) relies on the interreducibility of counting and sampling, and computes, for each pair of state $q$ and natural number $i\leq n$, a set of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ many uniformly chosen samples from the set of words of length $i$ that have a run ending at $q$. This informative measure — the number of samples maintained per state and length — also affects the overall time complexity with a quadratic dependence. Given the prohibitively high time complexity, in terms of each of the input parameters, of the FPRAS due to (Arenas et al., 2019), and considering the widespread application of approximate counting (and sampling) in various tasks in Computer Science, a natural question arises: Is there a faster FPRAS for #NFA that can pave the way for the practical implementation of approximate #NFA tools? In this work, we demonstrate that significant improvements in time complexity are achievable. Specifically, we have reduced the number of samples required for each state to be independent of $m$, with significantly less dependence on $n$ and $\epsilon$, maintaining only $\widetilde{O}(\frac{n^{4}}{\epsilon^{2}})$ samples per state. We believe our insights may lead to further reductions in time complexity and thus open up a promising avenue for future work towards the practical implementation of tools for approximating #NFA.
The authors decided to forgo the old convention of alphabetical ordering of authors in favor of a randomized ordering, denoted by r⃝. The publicly verifiable record of the randomization is available at https://www.aeaweb.org/journals/policies/random-author-order/search
copyright: noneconference: ; ;
@@ 1. Introduction
The counting problem #NFA and the closely related problem of uniform sampling words of an NFA are fundamental to database applications, and arise when counting and uniformly generating solutions to queries. Beyond database systems, the aforementioned problem finds applications in program verification (Legay et al., 2010), testing of software (Sutton et al., 2007) and distributed systems (Ozkan et al., 2019), probabilistic programming, estimation of information leakage in software (Bang et al., 2016; Gao et al., 2019; Saha et al., 2023), and even music generation (Donzé et al., 2014).
How do we count the number of elements of $\mathcal{L}(\mathcal{A}_{n})$? When $\mathcal{A}$ is a DFA with $m$ states, this question can be solved using a simple dynamic programming algorithm that inductively computes, for every $i\leq n$ and for every state $q$ of $\mathcal{A}$, the number of words of length $i$ that end in state $q$: $\mathsf{count}(q,i)=\sum_{b\in\Sigma}\sum_{p:q=\Delta(p,b)}\mathsf{count}(p,i-1)$ taking overall $O(mn)$ time (assuming $|\Sigma|=O(1)$). This simple algorithm however fails to work when $\mathcal{A}$ is a non-deterministic automaton as the same word maybe accepted using many different runs. Determinizing $\mathcal{A}$ first and then using the same algorithm as for DFAs would give an overall running time of $O(n\cdot 2^{m})$ which is exponential in the size of the description of the input. It turns out that #NFA is, in fact, provably computationally intractable — it is known to be a #P-hard problem (Álvarez and Jenner, 1993).
Given the intractability, and given the widespread applications of this problem, the next natural question then arises — can we approximate #NFA using a polynomial time algorithm? One of the promising paradigms for deriving efficient algorithms for approximating the answer to an intractable problem is the use of randomization to obtain a fully polynomial time randomized approximation scheme (FPRAS). Formally, an FPRAS for a real-valued function $f:\mathcal{D}\to\mathbb{N}$ over some domain $\mathcal{D}$ is a randomized algorithm $A:\mathcal{D}\times(0,1)\times(0,1)\to\mathbb{N}$ that for every input $x\in\mathcal{D}$ and parameters $\epsilon\in(0,1)$ and $\delta\in(0,1)$, runs in time $\textsf{poly}(|x|,\frac{1}{\epsilon},\log(\frac{1}{\delta}))$, and satisfies: ${\mathsf{Pr}}\!\left[\frac{1}{(1+\epsilon)}\cdot f(x)\leq|A(x,\epsilon,\delta)|\leq(1+\epsilon)\cdot f(x)\right]\geq 1-\delta$.
@@
Until very recently, the only approximation algorithms known for #NFA had running times that were quasi-polynomial in the size of the inputs (Gore et al., 1997; Kannan et al., 1995). In a recent breakthrough by Arenas, Croquevielle, Jayram and Riveros (Arenas et al., 2019, 2021), #NFA was finally shown to admit an FPRAS. Even though the problem was shown to admit an FPRAS, the time complexity of the proposed FPRAS is prohibitively large — it takes time $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, where the notation $\widetilde{O}$ hides the polylogarithmic multiplicative factors in $m$ and $n$. To be applicable in a practical setting like estimating information flow in programs, a polynomial time approximation scheme whose time complexity varies as $O(|P|^{16})$, where $|P|$ is roughly the lines of code, is unlikely to scale even for moderately sized programs. Towards this, in this work, we pursue the goal of desigining a faster algorithm, and propose an FPRAS which runs in time $\widetilde{O}(m^{4}n^{11}\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$. The reliance of both FPRASes on the inter-reducibility of sampling and counting necessitates a more informative measure, such as the number of samples maintained for each state. The scheme proposed in  (Arenas et al., 2021) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per state while the scheme proposed in this work maintains only $O(\frac{n^{4}}{\epsilon^{2}})$ samples, which is independent of $m$ and with a significant reduction in it’s dependence on $n$ and $\epsilon$.
@@ 1.1. Overview of our FPRAS
At a high-level, our approach exploits the tight connection between approximate counting and the related problem of (almost) uniform generation. In the context of an NFA $\mathcal{A}$ and a length $n$, the almost uniform generation problem asks for a randomized algorithm (generator) $G$, such that, on each invocation, the probability with which $G$ returns any given string $x\in\mathcal{L}(\mathcal{A}_{n})$ lies in the range $[\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}-\epsilon,\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}+\epsilon]$; here $\epsilon\in(0,1)$ is a parameter. Such a random generator is a polynomial time generator if its running time is polynomial in $|\mathcal{A}|$ and $n$. The existence of a polynomial time almost uniform generator for self-reducible (Jerrum et al., 1986) problems implies the existence of an FPRAS, and vice versa.
Input : NFA $\mathcal{A}$ with $m$ states $Q$, $n\in\mathbb{N}$ (in unary)
@@
2    Compute the estimate $\mathsf{N}({q}^{\ell})$ using prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$ Compute a uniform set of samples $\mathsf{S}({q}^{\ell})$ using $\mathsf{N}({q}^{\ell})$ and the prior estimates and samples $\{\mathsf{N}({q}^{i}),\mathsf{S}({q}^{i})\}_{i<\ell,p\in Q}$
return $\mathsf{N}({F}^{\ell})$
Algorithm 1 Algorithm Template: FPRAS for #NFA
This tight connection between counting and sampling was also the high-level idea behind the work of (Arenas et al., 2021). Consequently, our algorithm fits the template procedure of Algorithm 1, also outlined in (Arenas et al., 2021). For technical convenience, the algorithm unrolls the automaton $\mathcal{A}$ into an acyclic graph with $n$ levels; each level $\ell$ containing the $\ell^{\text{th}}$ copy ${q}^{\ell}$ of each state $q$ of $\mathcal{A}$. A key component in the algorithm is to incrementally compute an estimate $\mathsf{N}({q}^{\ell})$ of the size of the set $\mathcal{L}({q}^{\ell})$ of words of length $\ell$ that have a run ending in state $q$.
Next, in order to inductively estimate $\mathsf{N}({q}^{\ell})$, we in turn rely on the estimates of sets $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ (${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma=\{0,1\}$); here $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is a ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor state of $q$, i.e., $(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta$ is a transition of $\mathcal{A}$. These estimates can be put to use effectively because:
$\mathcal{L}({q}^{\ell})=\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\cup\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$

### 08 Apr 2024 v2

@@ Abstract.
Given a non-deterministic finite automaton (NFA) $\mathcal{A}$ with $m$ states, and a natural number $n\in\mathbb{N}$ (presented in unary), the #NFA problem asks to determine the size of the set $\mathcal{L}(\mathcal{A}_{n})$ of words of length $n$ accepted by $\mathcal{A}$. While the corresponding decision problem of checking the emptiness of $\mathcal{L}(\mathcal{A}_{n})$ is solvable in polynomial time, the #NFA problem is known to be #P-hard. Recently, the long-standing open question — whether there is an FPRAS (fully polynomial time randomized approximation scheme) for #NFA — was resolved by Arenas, Croquevielle, Jayaram, and Riveros in (Arenas et al., 2019). The authors demonstrated the existence of a fully polynomial randomized approximation scheme with a time complexity of $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, for a given tolerance $\epsilon$ and confidence parameter $\delta$. Given the prohibitively high time complexity in terms of each of the input parameters, and considering the widespread application of approximate counting (and sampling) in various tasks in Computer Science, a natural question arises: is there a faster FPRAS for #NFA that can pave the way for the practical implementation of approximate #NFA tools? In this work, we answer this question in the positive. We demonstrate that significant improvements in time complexity are achievable, and propose an FPRAS for #NFA that is more efficient in terms of both time and sample complexity.
A key ingredient in the FPRAS due to Arenas, Croquevielle, Jayaram, and Riveros (Arenas et al., 2019) is inter-reducibility of sampling and counting, which necessitates a closer look at the more informative measure — the number of samples maintained for each pair of state $q$ and length $i\leq n$. In particular, the scheme of (Arenas et al., 2019) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per pair of state and length. In the FPRAS we propose, we systematically reduce the number of samples required for each state to be only poly-logarithmically dependent on $m$, with significantly less dependence on $n$ and $\epsilon$, maintaining only $\widetilde{O}(\frac{n^{4}}{\epsilon^{2}})$ samples per state. Consequently, our FPRAS runs in time $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$. The FPRAS and its analysis use several novel insights. First, our FPRAS maintains a weaker invariant about the quality of the estimate of the number of samples for each state $q$ and length $i\leq n$. Second, our FPRAS only requires that the distribution of the samples maintained is close to uniform distribution only in total variation distance (instead of maximum norm). We believe our insights may lead to further reductions in time complexity and thus open up a promising avenue for future work towards the practical implementation of tools for approximate #NFA.
##\##NFA, counting, estimation, FPRAS
copyright: rightsretainedjournal: PACMMODjournalyear: 2024journalvolume: 2journalnumber: 2 (PODS)article: 112publicationmonth: 5doi: 10.1145/3651613ccs: Theory of computation Theory and algorithms for application domains
@@ 1. Introduction
The problem of #NFA is a fundamental problem in Computer Science, with range of applications in database and information extraction systems; we outline a few below.
Probabilistic Query Evaluation. Given a query $Q$ and a database $D$, the problem of probabilistic query evaluation (PQE) is to determine the probability of the query holding on a randomly sampled database, in which each row is included independently with probability given by its annotated value. PQE is known to be #P-hard, even in data complexity setting. For a large class of queries, PQE can directly be reduced to the #NFA problem. In particular, when the schema of the database $D$ consists only of binary relations, and when the query $Q$ is a self-join-free path query, then the PQE problem for $D$ and $Q$ reduces to an instance of #NFA, whose size is linear in the size of $Q$ and linear in the size of $D$ (van Bremen and Meel, 2023), and further, the time to compute the reduction is also linear in $D$ and $Q$. Thus an efficient algorithm for solving #NFA directly yields an efficient algorithm for probabilistic query evaluation.
Counting Answers to Reglar Path Queries. Regular path queries (or property paths(Angles et al., 2017) form a rich fragment of graph query languages such as SPARQL designed for graph database systems. Here, one abstracts information on edges in the graph database using an alphabet $\Sigma$ and asks a variety of questions about the set of paths that start from a given source node $u$, end at a target node $v$. In particular, a regular path query $(uRv)$ asks to enumerate, count or uniformly sample the set of paths (bounded in length by some fixed number $n$) from $u$ to $v$ whose labels match the regular expression $R$. Therefore, the problem of counting the number of answers to the queries reduces to the #NFA problem, for the NFA obtained by the cross product of NFA represented by the database (with $u$ as the intial state and $v$ as the final state) and the NFA that $R$ gets compiled down to. This reduced instance is again linear in the size of the database as well as the query, and the reduction takes the same time (Arenas et al., 2019).
Probabilistic Graph Homomorphism. A probabilistic graph is a pair $(H,\pi)$, where $H$ is a graph and $\pi$ is the associated probability labeling over edges. The associated probability space is defined as the set of all subgraphs of $H$ wherein every edge $e$ is sampled independently with probability $\pi(e)$. Given a query Graph $G$ and a probabilistic graph $(H,\pi)$ the problem of probabilistic graph homomorphism problem asks to to compute the probability that a randomly sampled subgraph of $H$ admits a homomorphism from G. In case of 1-way (and 2-way) path queries, the problem was again shown to reduce to #NFA (Amarilli et al., 2024).
@@
Beyond databases. From a complexity theoretic viewpoint, #NFA is SpanL-complete and therefore, finds applications in domains beyond databases such as program verification (Legay et al., 2010), testing of software (Sutton et al., 2007) and distributed systems (Ozkan et al., 2019), probabilistic programming, estimation of information leakage in software (Bang et al., 2016; Gao et al., 2019; Saha et al., 2023), information extraction (Arenas et al., 2021), automated reasoning using BDDs (Arenas et al., 2021), and even music generation (Donzé et al., 2014).
The direct applications of the #NFA problem brings up the central computational question — how do we count the number of elements of $\mathcal{L}(\mathcal{A}_{n})$? It turns out that #NFA is, in fact, provably computationally intractable — it is known to be a #P-hard problem (Álvarez and Jenner, 1993). Given the intractability, and given the widespread applications of the #NFA problem, the next natural question then arises — can we approximate #NFA using a polynomial time algorithm? Until very recently, the only approximation algorithms known for #NFA had running times that were quasi-polynomial in the size of the inputs (Gore et al., 1997; Kannan et al., 1995). In a recent breakthrough by Arenas, Croquevielle, Jayaram and Riveros (Arenas et al., 2019, 2021), #NFA was finally shown to admit fully polynomial time randomized approximation scheme (FPRAS).
One would assume that the discovery of FPRAS for #NFA would lead to design of scalable tools for all the different applications outlined above. Such has, however, not been the case and the primary hurdle to achieving practical adoption has been the prohibitively large time complexity of the proposed FPRAS– it takes time $\widetilde{O}\left(m^{17}n^{17}\cdot\frac{1}{\epsilon^{14}}\cdot\log(1/\delta)\right)$, where $\widetilde{O}$ hides the poly-logarithmic multiplicative factors in $m$ and $n$. It is worth highlighting that for most applications, the reduction to #NFA instance often only incurs a blow up that is a low-degree polynomial (linear or quadratic) factor larger than the size of the original problem. This means that the bottleneck to solving the original counting problem is really the high runtime complexity involved in the (approximate) counting algorithm and not the reduction.
@@
In this work, we pursue the goal of designing a faster algorithm, and propose an FPRAS which runs in time $\widetilde{O}((m^{2}n^{10}+m^{3}n^{6})\cdot\frac{1}{\epsilon^{4}}\cdot\log^{2}(1/\delta))$. The reliance of both FPRASes on the inter-reducibility of sampling and counting necessitates a more informative measure, such as the number of samples maintained for each state. The scheme proposed in  (Arenas et al., 2021) maintains $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ samples per state while the scheme proposed in this work maintains only $O(\frac{n^{4}}{\epsilon^{2}})$ samples, which is independent of $m$ and with a significant reduction in its dependence on $n$ and $\epsilon$.
@@ 1.1. Overview of our FPRAS
At a high-level, our approach exploits the tight connection between approximate counting and the related problem of (almost) uniform generation. In the context of an NFA $\mathcal{A}$ and a length $n$, the almost uniform generation problem asks for a randomized algorithm (generator) $G$, such that, on each invocation, the probability with which $G$ returns any given string $x\in\mathcal{L}(\mathcal{A}_{n})$ lies in the range $[\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}(1-\epsilon),\frac{1}{|\mathcal{L}(\mathcal{A}_{n})|}(1+\epsilon)]$; here $\epsilon\in(0,1)$ is a parameter. Such a random generator is a polynomial time generator if its running time is polynomial in $|\mathcal{A}|$, $n$ and $1/\epsilon$. The existence of a polynomial time almost uniform generator for self-reducible (Jerrum et al., 1986) problems implies the existence of an FPRAS, and vice versa.
This tight connection between counting and sampling was also the high-level idea behind the work of (Arenas et al., 2021). Consequently, our algorithm fits the template procedure of Fig. 1, also outlined in (Arenas et al., 2021). For technical convenience, the algorithm unrolls the automaton $\mathcal{A}$ into an acyclic graph [^1] The assumption in the template algorithm of Fig. 1 that $\mathcal{A}$ has only 1 accepting state is without loss of generality. with $n$ levels; each level $\ell$ containing the $\ell^{\text{th}}$ copy ${q}^{\ell}$ of each state $q$ of $\mathcal{A}$. A key component in the algorithm is to incrementally compute an estimate $\mathsf{N}({q}^{\ell})$ of the size of the set $\mathcal{L}({q}^{\ell})$ of words of length $\ell$ that have a run ending in state $q$. Next, in order to inductively estimate $\mathsf{N}({q}^{\ell})$, we in turn rely on the estimates of sets $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})$ and $\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})$; here $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is a ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor state of $q$ (${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma=\{0,1\}$), i.e., $(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta$ is a transition of $\mathcal{A}$. These estimates can be put to use effectively because:
$\mathcal{L}({q}^{\ell})=\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}\}\uplus\big{(}\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}}}^{\ell-1})\big{)}\cdot\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}.$
Figure 1. Algorithm Template: FPRAS for #NFA
Input : NFA $\mathcal{A}$ with $m$ states $Q$ and a single final state $q_{F}\in Q$, $n\in\mathbb{N}$ (in unary)
@@
2    Compute the estimate $\mathsf{N}({q}^{\ell})$ using prior estimates and samples $\{\mathsf{N}({p}^{i}),\mathsf{S}({p}^{i})\}_{i<\ell,p\in Q}$ Compute a uniform set of samples $\mathsf{S}({q}^{\ell})$ using $\mathsf{N}({q}^{\ell})$ and the prior estimates and samples $\{\mathsf{N}({p}^{i}),\mathsf{S}({p}^{i})\}_{i<\ell,p\in Q}$
return $\mathsf{N}({q_{F}}^{n})$